How much would you age in traveling a distance of 4.66 light years from earth if you were able to travel at a speed of 0.763c?

I got 6.11 years but this is not correct.
Answer #1

first, we will find time decay constant

= 1 / sqrt (1 - 0.7632 c2 / c2)

= 1.547

time passed on earth = 4.66 / 0.763 = 6.107 years

so,

time passed (age) = 6.107 / 1.547

age = 3.947 years